The main reason that distillation is a costly method for purifying water is the high energy required to heat and vaporize water.

Using the density, specific heat, and heat of vaporization of water from Appendix B in the textbook, calculate the amount of energy required to vaporize 1.50 gal of water beginning with water at 24 ∘C.

Express your answer using two significant figures.

E =   kJ  

SubmitMy AnswersGive Up

Part B

If the energy is provided by electricity costing $9.3×10−2 /kwh, calculate its cost.

Express your answer using two significant figures.

$

SubmitMy AnswersGive Up

Part C

If distilled water sells in a grocery store for $1.25 per gal, what percentage of the sales price is represented by the cost of the energy?

Express your answer using two significant figures.

  %  
Answer #1

part A:

Q absorbed = m*s*DT + n*DHvap    (1.5 gal = 6607.3257 grams)

         = 6607.3257*4.18*(100-24) + (6607.3257/18)*(40.7*10^3)


   = 17038.9 kJ

part B :

1kwh = 10^3*3600 = 3.6*10^3 kJ

nO Of kwh = 17038.9/3600 = 4.733 kwh

cost of power = 4.733*0.093 = 0.4402 dolar

part C

cost of distilled = 1.5*1.25 = 1.875 dolar

percentage = 0.4402/1.875*100 = 23.48%